# Archetype-Aware Predictive Autoscaling with Uncertainty Quantification for Serverless Workloads on Kubernetes

### v1

@@ IV-C Baselines
We compare AAPA against two baselines:
Kubernetes HPA: Standard reactive autoscaler with 70% CPU target, 5-minute stabilization window, and scale-down cooldown of 5 minutes.
Generic Predictive: Applies Holt-Winters forecasting uniformly across all workloads with 15-minute prediction horizon [^1] Implementation based on https://github.com/jthomperoo/predictive-horizontal-pod-autoscaler.
Our baseline selection aligns with recent systematic evaluations of ML-based autoscalers, which identified prediction accuracy-latency tradeoffs across different regression models [16]. We chose these baselines to represent both reactive (HPA) and predictive approaches without workload differentiation.
@@ IV-D Evaluation Metrics
We evaluate autoscalers across three dimensions:
Performance metrics:
  • SLO violation rate: percentage of requests exceeding 500ms response time
  • Cold start frequency: requests arriving to zero available pods
  • P95/P99 response times
Efficiency metrics:
  • Total replica-minutes: integral of active replicas over time
  • Average CPU utilization
  • Over-provisioning rate: time with utilization ¡50%
Stability metrics:
  • Scaling oscillations: consecutive scale-up/down actions
  • Average interval between scaling decisions
IV-E Experimental Setup
Each experiment simulates one day of workload replay with:
  • Initial replica count: 2
  • Maximum replicas: 100
  • CPU capacity: 1000 millicores per replica
  • Memory: 256MB per replica (from dataset averages)
We run 5 trials with different random seeds and report averaged results with 95% confidence intervals. All experiments use the same workload traces to ensure fair comparison.
@@ V Results and Analysis
We evaluated AAPA against baseline autoscalers using 300,000 sliding window samples extracted from the Azure Functions 2019 dataset. Our experiments reveal important insights about the performance-cost tradeoffs in ML-driven autoscaling.
@@ V-A Workload Classification Performance
Our weak supervision approach successfully labeled the dataset with high accuracy. The LightGBM classifier achieved 99.8% accuracy on the test set. To handle class imbalance (PERIODIC: 70.2%, SPIKE: 17.6%, STATIONARY_NOISY: 12.0%, RAMP: 0.2%), we applied class weights inversely proportional to frequency during training.
@@
True/Pred PERI SPIKE STAT RAMP
@@
Our experiments reveal a fundamental tradeoff between performance optimization and resource efficiency. Figure 2 presents a multi-dimensional analysis of this tradeoff.
@@ V-B2 Resource Utilization Analysis
The cost of AAPA’s performance improvements is substantial:
@@
This stems from warm pool maintenance (1-3 additional pods), conservative utilization targets (30-65%), and predictive over-provisioning.
@@ V-C Perspective-Aware Discussion
For SPIKE workloads: AAPA’s 7.7× resource overhead provides warm pools that eliminate cold starts and handle bursts effectively. However, the 5-second scaling interval limits response to extremely short spikes. Cloud providers may find this cost prohibitive, while latency-sensitive applications may justify the expense.
For PERIODIC workloads: All autoscalers achieve ¿98% SLO compliance, making AAPA’s 2× resource overhead difficult to justify. Simple reactive scaling suffices when workloads are predictable.
@@ V-D Statistical Significance and Variability
Wilcoxon signed-rank tests ($\alpha=0.05$) confirm statistical significance (p¡0.01) for all performance improvements. AAPA shows higher resource variability than HPA, indicating workload sensitivity.
Sensitivity analysis on REI weights shows that varying $\alpha$, $\beta$, $\gamma$ by ±$\pm 0.05$ changes autoscaler rankings by less than 2%, confirming robustness of our conclusions.
V-E Key Findings and Implications
Key findings: (1) No universal winner—optimal autoscaler depends on priorities; (2) 99.8% accurate classification enables dynamic strategy selection; (3) Warm pools eliminate cold starts despite resource costs; (4) ML shifts but cannot eliminate fundamental cost-performance tradeoffs. AAPA’s value lies in providing predictable performance when it matters most.
@@ VI-B Practical Deployment Considerations
Industrial best practices [17] recommend starting with reactive scaling (HPA/VPA) and progressively adopting ML-based approaches for workloads exhibiting clear patterns. AAPA’s archetype-based design aligns with this philosophy by automatically identifying which workloads benefit from sophisticated strategies. HPEC facilities should adopt hybrid approaches—AAPA for real-time data paths, HPA for offline analytics—while monitoring actual resource utilization to validate the cost-performance tradeoffs.
@@ VII Conclusion
This paper presented AAPA, an archetype-aware predictive autoscaler for serverless workloads on Kubernetes. We developed a weak supervision framework that automatically labeled 300,000 time windows into four workload archetypes with 99.8% accuracy, eliminating manual annotation. Our 37-feature engineering pipeline and differentiated scaling strategies demonstrate how ML can navigate, rather than eliminate, fundamental performance-cost tradeoffs in distributed systems.
Our experiments on Azure Functions traces reveal clear quantitative tradeoffs: AAPA reduces SLO violations by up to 50% and maintains sub-200ms response times for 95% of requests, but requires 2-8× more resources than traditional HPA. SPIKE workloads show the highest resource overhead (7.7×) but benefit most from warm pools and predictive scaling. PERIODIC workloads achieve ¿98% SLO compliance with all approaches, making AAPA’s 2× overhead difficult to justify. The 2-second pod startup time emerges as a fundamental constraint that necessitates proactive resource allocation.
We recommend deploying AAPA for revenue-critical, user-facing services where SLO violations directly impact business metrics—payment APIs, authentication services, and customer-facing endpoints justify the resource premium. Traditional HPA remains optimal for batch processing, internal microservices, and cost-sensitive workloads. Organizations should implement hybrid strategies: use AAPA during peak hours or for spike-prone services, while applying HPA elsewhere. Our 99.8% accurate classifier enables dynamic strategy selection based on workload characteristics.
Future work should explore online parameter adaptation to reduce resource waste and multi-objective optimization with explicit cost budgets. Our work ultimately reframes ML’s role in systems: rather than magical optimization, ML provides intelligent navigation of fundamental tradeoffs. As serverless adoption grows in HPEC environments, automated approaches like our weak supervision framework become essential for managing operational complexity at scale. We invite the community to reproduce and extend our results using the provided traces and simulator at https://github.com/GuilinDev/aapa-simulator.
@@ Acknowledgments
The authors thank the DPOE-Scout and DPOE-Insights teams at Workday Inc. for their valuable insights on industrial workload patterns and helpful discussions that validated our approach. This research uses only publicly available datasets; no Workday proprietary data was used in this study.

### v3

@@ IV-C Baselines
We compare AAPA against two widely adopted autoscaling strategies:
  • Kubernetes HPA: The default reactive policy using a 70% CPU target, 5-minute stabilization window, and scale-down cooldown.
  • Generic Predictive Autoscaler: A uniform Holt-Winters forecasting model with a 15-minute prediction horizon applied across all workloads. [^2] https://github.com/jthomperoo/predictive-horizontal-pod-autoscaler
These baselines represent common reactive and predictive paradigms, both of which lack workload differentiation. Our selection aligns with recent evaluations of machine learning–based autoscalers [24].
@@ IV-D Evaluation Metrics
We use fine-grained metrics to assess autoscaling behavior from three perspectives:
  • Performance: SLO violation rate (requests exceeding 500ms), cold start frequency, and P95/P99 response latency.
  • Efficiency: Total replica-minutes (area under the replica curve), average CPU utilization, and underutilization rate (fraction of time with CPU $<$50%).
  • Stability: Number of scaling oscillations and average time between scaling events.
These metrics also serve as building blocks for the REI metric (Section III-C), which aggregates them into a single score for high-level comparison.
@@ V Results and Analysis
We now present key findings based on the experimental setup described in Section IV, focusing on the performance–cost tradeoffs observed across diverse workload archetypes.
@@ V-A Workload Classification Performance
Our weak supervision approach successfully labeled the dataset with high fidelity. The LightGBM classifier achieved 99.8% accuracy on the test set. While the overall dataset exhibits class imbalance (PERIODIC: 70.2%, SPIKE: 17.6%, STATIONARY_NOISY: 12.0%, RAMP: 0.2%), the test set distribution differs as noted in Section IV.
The confusion matrix in Table IV shows near-perfect classification across all archetypes, including the minority RAMP class, which achieves 95.0% precision and 96.6% recall despite its scarcity.
@@
True/Pred PERIODIC SPIKE STAT. RAMP
@@
Our experiments reveal a fundamental tradeoff between performance optimization and resource efficiency. Fig. 2 presents a multi-dimensional analysis of this tradeoff.
@@ V-B2 Resource Utilization Analysis
The performance improvements achieved by AAPA come at a substantial cost, as shown in Table V. This overhead primarily stems from warm pool maintenance (1–3 additional pods), conservative utilization targets (30–75%), and predictive over-provisioning.
@@ V-C Perspective-Aware Discussion
For SPIKE workloads: AAPA’s 7.7× resource overhead (Table V) provides warm pools that eliminate cold starts and handle bursts effectively. However, the 5-second scaling interval limits response to extremely short spikes. Cloud providers may find this cost prohibitive, while latency-sensitive applications may justify the expense.
For PERIODIC workloads: All autoscalers achieve $>$98% SLO compliance, making AAPA’s 2× resource overhead (Table V) difficult to justify. Simple reactive scaling suffices when workloads are predictable. This is confirmed by our statistical tests showing no significant performance differences (p=0.621 for SLO violations, p=0.892 for response time).
@@ V-D Statistical Significance and Variability
Table VI presents our statistical validation results. Wilcoxon signed-rank tests ($\alpha=0.05$) confirm statistical significance for key performance improvements, particularly for SPIKE and STATIONARY workloads where AAPA’s adaptive strategies provide the greatest benefit. While AAPA shows higher resource variability than HPA (indicating workload sensitivity), our sensitivity analysis demonstrates that varying REI weights by ±$\pm 0.05$ changes autoscaler rankings by less than 2%, confirming the robustness of our conclusions.
TABLE VI: Statistical Validation of Results
Part A: Wilcoxon Signed-Rank Test Results (AAPA vs HPA)
Pattern SLO Violations Response Time Resource Usage
SPIKE 0.008** $<$0.001*** $<$0.001***
PERIODIC 0.621 0.892 $<$0.001***
RAMP 0.042* 0.156 $<$0.001***
STATIONARY 0.003** 0.009** $<$0.001***
Part B: REI Sensitivity Analysis
Weight Variation Original REI Adjusted REI Rank Change (%)
Baseline ($\alpha$=0.4, $\beta$=0.3, $\gamma$=0.3) HPA $>$ AAPA HPA $>$ AAPA 0.0
$\alpha$$\pm$ 0.05 HPA $>$ AAPA HPA $>$ AAPA 1.8
$\beta$$\pm$ 0.05 HPA $>$ AAPA HPA $>$ AAPA 1.2
$\gamma$$\pm$ 0.05 HPA $>$ AAPA HPA $>$ AAPA 0.6
Notes: * p$<$0.05, ** p$<$0.01, *** p$<$0.001. Lower p-values indicate stronger evidence against the null hypothesis of no difference between AAPA and HPA performance.
@@ VI-B Practical Deployment Considerations
Industrial best practices [25, 26] recommend starting with reactive scaling (HPA/VPA) and progressively adopting ML-based approaches for workloads exhibiting clear patterns. AAPA’s archetype-based design aligns with this philosophy by automatically identifying which workloads benefit from sophisticated strategies. HPEC facilities should adopt hybrid approaches—AAPA for real-time data paths, HPA for offline analytics—while monitoring actual resource utilization to validate the cost-performance tradeoffs.
@@ VII Conclusion
We presented AAPA, an archetype-aware autoscaler for serverless workloads, guided by a 99.8%-accurate classifier trained on our weakly labeled dataset AAPAset. AAPA applies archetype-specific, uncertainty-aware scaling strategies that reduce SLO violations by up to 50%, albeit with 2–8$\times$ higher resource costs under bursty conditions. Our findings reaffirm that ML shifts rather than eliminates the cost-performance tradeoff in autoscaling. AAPA is best suited for latency-critical, user-facing services, while traditional HPA remains sufficient for predictable or cost-sensitive workloads. We advocate hybrid deployments and dynamic strategy selection. By releasing AAPAset and proposing the REI metric, we provide practical tools for reproducible, workload-aware autoscaler evaluation. Future work will explore online adaptation and multi-objective scaling under cost constraints.
@@ Acknowledgments
The authors thank the DPOE-Scout and DPOE-Insights teams at Workday Inc. for insightful feedback on industrial workload patterns and for discussions that helped validate this work. This study relies exclusively on publicly available datasets; no proprietary Workday data were used, and the views expressed are solely those of the authors.